GE wants to do an investment in Brazil. The equity investment is BRL 100M and Debt issue is BRL 150 M. The analyst finds that Brazilian tax rate is 35% and the risk free rate is 7.4%. GE can borrow in Brazil at 60 bps over the Brazilian Treasuries. Similar projects in Brazil have a beta of 1.1. The return of the Brazilian market (BOVESPA) in the past 10 years is 12%. What would be the project’s cost of capital?
Answer #1

Equity investment:

Market value of equity, MV(Equity) = BRL 100m

  • risk-free rate = 7.4%
  • Return on BOVESPA market, Rm = 12%
  • beta of the project = 1.1
  • cost of equity using CAPM, Re = Rf + beta * (Rm - Rf)
  • Re = 7.4% + [ 1.1 * (12% - 7.4%)
  • Re = 7.4% + [ 1.1 * 4.60% ]
  • Re = 7.4% + 5.06% = 12.46%

Debt Investment:

Market value of debt, MV(Debt) = 150m

tax rate = 35%

Borrowing rate, cost of debt, Rd= 60bps over Brazilian treasuries = 60bps + 7.40% = 0.60% + 7.40% = 8%

Cost of Capital:

wacc = WtE * Re + WtD * (1-tax) * Rd

  • Equity = 100m
  • Debt = 150m
  • Capital = Debt + Equity = 100m + 150m = 250m
  • WtE = 100m / 250m = 0.40
  • WtD = 150m/250m = 0.60
  • wacc = 0.40 * 12.46% + 0.06 * ( 1 - 0.35) * 8% = 0.04984 + 0.0312 = 0.08104 = 8.104%

cost of capital = 8.104% (answer)